According to Chebyshev's theorem, what proportion of a distribution will be within k = 4 standard deviations of the mean? Show all work as to how to find this. please work the problem out. Thanks
Answer by rothauserc(4718)

note that 4^2 = 16
For K = 4 we have 1 – 1/K^2 = 1 - 1/16 = 15/16 = 93.75%. So Chebyshev’s theorem (inequality) says that at least 93.75% of the data values of any distribution must be within 4 standard deviations of the mean.